# SiniticMTError: A Machine Translation Dataset with Error Annotations for Sinitic Languages

### v1

@@ Abstract
Despite major advances in machine translation (MT) in recent years, progress remains limited for many low-resource languages that lack large-scale training data and linguistic resources. Cantonese and Wu Chinese are two Sinitic examples, although each enjoys more than 80 million speakers around the world. In this paper, we introduce SiniticMTError, a novel dataset that builds on existing parallel corpora to provide error span, error type, and error severity annotations in machine-translated examples from English to Mandarin, Cantonese, and Wu Chinese. Our dataset serves as a resource for the MT community to utilize in fine-tuning models with error detection capabilities, supporting research on translation quality estimation, error-aware generation, and low-resource language evaluation. We report our rigorous annotation process by native speakers, with analyses on inter-annotator agreement, iterative feedback, and patterns in error type and severity.
@@
Hannah Liu1, Junghyun Min2, Ethan Yue Heng Cheung1, Shou-Yi Hung1, Syed Mekael Wasti3, Runtong Liang1, Shiyao Qian1, Shizhao Zheng 1, Elsie Chan1, Ka Ieng Charlotte Lo1, Wing Yu Yip1, Richard Tzong-Han Tsai4, En-Shiun Annie Lee1,3 1University of Toronto   2Georgetown University 3Ontario Tech University   4National Central University, Taiwan hannahhere.liu@mail.utoronto.ca
1 Introduction
Machine translation (MT) has made significant advancements in recent years as either supervised systems (Luong et al., 2015; Lakew et al., 2018; Liu et al., 2020; Wang et al., 2022; Liu, 2022; Park et al., 2023) or LLMs Zhu et al. (2024); Freitag et al. (2024). However, such work often focuses on higher-resource languages, and progress remains limited with low-resource languages (Ranathunga et al., 2023), where fine-tuned models suffer from poor performance (Lee et al., 2022; Shliazhko et al., 2024) and LLMs output noise Iyer et al. (2024); Levine et al. (2025).
We focus on two major yet low-resource Sinitic languages, Cantonese [^1] Also known as Yue (Eberhard et al., 2023) and Wu Chinese [^2] Whose most well-known dialect is Shanghainese (Eberhard et al., 2023). We discuss our terminology for these languages in Appendix A.. They remain underserved despite having more than 80 million and 83 million speakers respectively, across southern and eastern China and various diasporas across the world Chappell (2015); Eberhard et al. (2023). The limited progress is attributable to the dominance of Mandarin as lingua franca in these regions Norman (1988); Li (2006), scarcity in publicly available parallel corpora Xiang et al. (2024), the lack of standardization in writing systems Pan et al. (1991); Tang et al. (2002); Kwan-hin and Bauer (2002); Snow (2008), and their status as primarily vernacular (i.e. spoken rather than written) languages Pan et al. (1991); Snow (2004); Li (2006).
Figure 1: Sample Mandarin entry. mt looks fluent, but contains subtle semantic errors: an unwarranted subjective phrase (Addition) and a lexical mistranslation (Mistranslation). While 漂亮 directly translates to beautiful, it usually describes people or objects and 好 good is more natural when used to describe the weather.
@@ 1 Introduction
In this paper, we add to the Sinitic MT literature by presenting SiniticMTError, a novel suite of datasets that build on FLORES+ (Goyal et al., 2022; NLLB Team et al., 2024; Yu et al., 2024) to provide erroneous machine translation examples and detailed span-level error annotations including error type and severity for Mandarin, Cantonese, and Wu Chinese (Figure 1).
We report preliminary statistics across 2,000 Mandarin and 1,000 Cantonese sentences, with Wu Chinese annotations still in progress. Early results suggest distinctive error distributions across languages. We describe our annotation process conducted by native speakers (Figure 2; Section 3), discussions on patterns in error types and severity (Section 4), and detailed analyses on translation issues (Section 5), offering both linguistic insights and resources to advance low-resource MT. A sample of our dataset is available at an Anonymous GitHub Repository.
2 Related Work
@@
Annotation frameworks like Multidimensional Quality Metrics (MQM; Burchardt, 2013) and Error Span Annotations (ESA; Chen et al., 2020; Kocmi et al., 2024, inter alia) represent important steps toward more equitable multilingual NLP by introducing standardized methods for evaluating translation quality. MQM and ESA facilitate identifying and localizing errors to improve translation pipelines Chen et al. (2020); Zhang et al. (2025), and explicit fine-tuning or prompting to improve the quality of generative model output Kocmi and Federmann (2023); Wang et al. (2024).
@@
While these frameworks offer a promising foundation, existing datasets built on them have largely focused on high-resource languages such as English, French, and Mandarin Chinese Freitag et al. (2021); Sellam et al. (2021). One effort to adapt these tools to lower-resource languages is by Singh et al. (2024) who focus on low-resource Indic languages.
@@
FLORES-101, FLORES-200, and FLORES+ Goyal et al. (2022); NLLB Team et al. (2024) are initiatives that have introduced multilingual benchmarks covering over 100 languages, many of which are low-resource, to support equitable evaluation of MT systems. Building upon FLORES-200, AfriCOMET Wang et al. (2024) is an annotation effort aiming to bridge the gap between low-resource languages and MT systems by providing annotated datasets for underrepresented African languages. Similar work has emerged for Bambara (Dou and Neubig, 2022), Amharic and Tigrinya Shapiro et al. (2023), and Spanish languages (Perez-Ortiz et al., 2024), offering parallel corpora and baseline models. Within the Sinitic language family, resources in Hokkien [^3] Also known as Min Nan (Southern Min; Eberhard et al., 2023) (Lu et al., 2022; Chen et al., 2023; Lu et al., 2024) and Hakka (Hung and Huang, 2022; Lai et al., 2024) have been compiled.
@@
Despite related efforts, publicly available resources in Cantonese and Wu Chinese remain scarce. Proprietary LLMs offer commercial service in Cantonese OpenAI et al. (2024); Team et al. (2024) and several Cantonese–Mandarin or Cantonese–English parallel corpora exist in the form of subtitles Wong and Zhang (2017), dictionaries Mair and DeFrancis (2003), or government transcripts Lee (2011). However, resources are limited in scale or accessibility as surveyed by Xiang et al. (2024). As a result, previous work in Cantonese MT have relied on synthetic data augmentation (Liu, 2022; Hong et al., 2024). To the best of our knowledge, the recent addition to FLORES+ Yu et al. (2024) represents the sole publicly available MT resource in Wu Chinese.
3 Dataset Construction and Annotation
@@
Each example in the dataset consists of a triplet of sentences: a source sentence (src), a machine-translated sentence (mt), and a reference translation (ref). The src and ref pairs are drawn from the FLORES+, which is an extension of the FLORES-200 dataset (NLLB Team et al., 2024). The mt sentences are generated using the 600M NLLB-200 Team et al. (2022) for Mandarin and Cantonese, and Qwen2.5 Max (Team, 2024) for Wu Chinese. We describe our model selection and mt generation in detail in Appendix B.
Figure 3: Annotators were presented with the original English source sentence, the machine-translated output (mt), and a human reference translation for comparison. The green check marks indicate completed annotations.
@@
The constructed dataset is then provided to annotators for error span annotation. For each language pair (English-Mandarin, English-Cantonese, English-Wu Chinese), we recruit three bilingual annotators who are listed as collaborating authors. The annotators are native speakers of their respective target languages and highly proficient in English. They are familiar with both the linguistic conventions of their target language and the goals of the annotation task.
We use the TranslationCorrect tool (Wasti et al., 2025) as our annotation interface. A sample screen from the interface is illustrated in Figure 3. We discuss the interface in greater detail in Appendix C.
3.1 Annotation Guidelines
@@
Annotators are instructed to examine the mt sentence with reference to both src and ref, and identify any translation errors by highlighting spans directly on the mt sentences. For each error span, annotators categorize the severity and error type, while recording the erroneous span indices in the mt sentence. The label spaces for severity and type, adopted from MQM guidelines (Burchardt, 2013) and the AfriCOMET framework (Wang et al., 2024) are described in Appendix D.
@@
Granularity. When identifying errors, the annotators are asked to be as fine-grained as possible. For example, if a sentence contains two words that are each mistranslated, two separate mistranslation error spans should be recorded. If a single text segment contains multiple errors, the leftmost span with highest severity is to be recorded.
@@
Additional info. One error type is omission, where content in the source sentence is missing from the translation. In such case, annotators are instructed to insert the missing information the post-editing box and highlight the inserted span with the appropriate type (omission), span, and severity labels.
3.2 Annotation Workflow
@@
Mandarin pilot. Mandarin annotators had two rounds of pilot studies. In the first round, each annotator completed 50 examples, which were then reviewed by a language lead. The lead provided feedback and held group discussions to ensure consistency, before annotators completed a second round of 50 examples. Qualitative analysis showed clear improvement in annotation consistency.
Cantonese pilot. Cantonese annotators followed a more granular setup consisting of 4-rounds. The initial round included 50 examples and the others 10 examples each, completing 80 examples in total. Additional rounds of 10 examples were added to ensure sufficient agreement before proceeding. We describe the details of adjustments made during the pilot stages in Appendix E.
Main annotation phase. During the main annotation phase, annotators worked in batches of 50 to 200 sentences. With each batch, we employed a similar iterative process where annotations were reviewed, recurring issues discussed, and relevant guidelines fine-tuned. During such iterative process, guidelines on annotation granularity and categorizing missing punctuation as omission were established.
Quality assurance. After completing the first round of annotations (main annotator round), annotators proceed to the Quality Assurance (QA) stage. In this phase, each annotator will be assigned sentences previously annotated by another team member. Each participant reviews these annotations and discusses any discrepancies with the original annotator, should there be any, to improve consistency across annotators. We plan to perform inter-annotator agreement analysis in multiple stages.
4 Preliminary Dataset Statistics
We report our preliminary statistics from the ongoing annotation of Mandarin and Cantonese data. At the time of writing, we have annotated 2009 Mandarin sentences and 996 Cantonese sentences, with Wu Chinese still in its pilot training phase. Table 1 and Table 2 present the distribution of error categories and severity levels in our current annotation data. On average, each Mandarin sentence contains 1.93 annotation spans, while each Cantonese sentence contains 6.82 spans, showing a much higher span density than Mandarin. For Wu Chinese, we will follow the same annotation setup as Mandarin.

### v2

@@ Abstract
Despite major advances in machine translation (MT) in recent years, progress remains limited for many low-resource languages that lack large-scale training data and linguistic resources. In this paper, we introduce SiniticMTError, a novel fine-grained dataset that builds on existing parallel corpora to provide error span, error type, and error severity annotations in machine-translated examples from English to Mandarin, Cantonese, and Wu Chinese, along with a Mandarin-Hokkien component derived from a non-parallel source. Our dataset serves as a resource for the MT community to fine-tune models with error detection capabilities, supporting research on translation quality estimation, error-aware generation, and low-resource language evaluation. We also establish baseline results using language models to benchmark translation error detection performance. Specifically, we evaluate multiple open source and closed source LLMs using span-level and correlation-based MQM metrics, revealing their limited precision, underscoring the need for our dataset. Finally, we report our rigorous annotation process by native speakers, with analyses on pilot studies, iterative feedback, insights, and patterns in error type and severity.
Keywords: machine translation, error annotation, Sinitic language
[…]
@@
Hannah Liu1, Junghyun Min2, En-Shiun Annie Lee1,3, Ethan Yue Heng Cheung1,
Shou-Yi Hung1, Elsie Chan1, Shiyao Qian1, Runtong Liang1, Kimlan Huynh1,
Wing Yu Yip1, York Hay Ng1, TSZ Fung Yau1, Ka Ieng Charlotte Lo1, You-Wei Wu4,
Richard Tzong-Han Tsai4
1University of Toronto, 2Georgetown University
3Ontario Tech University, 4National Central University, Taiwan
hannahhere.liu@mail.utoronto.ca
Abstract content
1.  Introduction
Machine translation (MT) systems have made significant advancements in recent years both through supervised systems (Luong et al., 2015; Lakew et al., 2018; Liu et al., 2020; Wang et al., 2022; Liu, 2022; Park et al., 2023) and through large language models (LLMs) Zhu et al. (2024); Freitag et al. (2024). However, such systems often focus on higher-resource languages, and progress remains limited with low-resource languages (Ranathunga et al., 2023), where fine-tuned models suffer from poor performance (Lee et al., 2022; Shliazhko et al., 2024) and LLMs output noise Iyer et al. (2024); Levine et al. (2025).
Figure 1: Sample Mandarin entry. mt looks fluent, but contains subtle semantic errors: an unwarranted subjective phrase (Addition) and a lexical mistranslation (Mistranslation). While 漂亮 piao4liang directly translates to beautiful, it usually describes people or objects and 好 good is more natural when used to describe the weather.
In addition to Mandarin, we focus on three major yet low-resource Sinitic variants, Cantonese [^1] Also known as Yue (Eberhard et al., 2023), Wu Chinese [^2] Whose most well-known dialect is Shanghainese (Eberhard et al., 2023), and Hokkien [^3] Also known as Min Nan (Southern Min; Eberhard et al., 2023). They remain underserved despite having more than 80, 83 and 47 million speakers respectively, across southern and eastern China and various diasporas across the world Chappell (2015); Eberhard et al. (2023). The limited progress is attributable to the dominance of Mandarin as lingua franca in these regions Norman (1988); Li (2006), scarcity in publicly available parallel corpora Xiang et al. (2024), the lack of standardization in writing systems Pan et al. (1991); Tang et al. (2002); Kwan-hin and Bauer (2002); Snow (2008), and their status as primarily vernacular (i.e. spoken rather than written) languages Pan et al. (1991); Snow (2004); Li (2006).
@@ 1.  Introduction
In this paper, we add to the Sinitic MT literature by presenting SiniticMTError, a novel fine-grained suite of datasets that mainly build on FLORES+ (Goyal et al., 2022; NLLB Team et al., 2024; Yu et al., 2024) to provide erroneous machine translation examples and detailed span-level error annotations including error type and severity for Mandarin, Cantonese, Wu Chinese, and Hokkien (Figure 1).
We report dataset statistics across 2,009 Mandarin and 1,402 Cantonese sentences, 68 Wu Chinese and 154 Hokkien pilot annotations. Our results suggest distinctive error distributions across languages. The currently available annotated portions of the dataset can be found in our GitHub Repository.
2.  Background
We discuss relevant annotation frameworks, resources for multilingual and Sinitic machine translation, and clarify the language terminology we use in this paper.
2.1.  Literature Review
Annotation frameworks.
@@ 2.1.  Literature Review
Annotation frameworks like Multidimensional Quality Metrics (MQM; Lommel et al., 2013) and Error Span Annotations (ESA; Chen et al., 2020; Kocmi et al., 2024, inter alia) represent important steps toward more equitable multilingual NLP by introducing standardized methods for evaluating translation quality. MQM and ESA facilitate identifying and localizing errors to improve translation pipelines Chen et al. (2020); Zhang et al. (2025a), and explicit fine-tuning or prompting to improve the quality of generative model output Kocmi and Federmann (2023). SiniticMTError annotations and design choices, including annotation instructions, error type and data structure are based on MQM (Lommel et al., 2013) and AfriCOMET, an adaptation of the MQM framework for African languages (Wang et al., 2024).
Multilingual resources.
@@
While MQM and ESA offer a promising foundation, existing datasets built on them have largely focused on high-resource languages such as English, French, and Mandarin Chinese Freitag et al. (2021); Sellam et al. (2021). Efforts to adapt these tools to lower-resource languages include Singh et al. (2024); Wang et al. (2024); Li et al. (2025), which focus on low-resource Indic and African languages.
@@
FLORES-101, FLORES-200, and FLORES+ Goyal et al. (2022); NLLB Team et al. (2024) are initiatives that have introduced multilingual benchmarks covering over 100 languages, many of which are low-resource, to support equitable evaluation of MT systems. Building upon the developmental and test splits of FLORES-200, AfriCOMET Wang et al. (2024) is an annotation effort aiming to bridge the gap between low-resource languages and MT systems by providing annotated datasets for underrepresented African languages. Similar work has emerged for Bambara (Dou and Neubig, 2022), Amharic and Tigrinya Shapiro et al. (2023), and Spanish languages (Perez-Ortiz et al., 2024), offering parallel corpora and baseline models.
Sinitic resources.
@@
Despite related efforts, publicly available resources in Cantonese, Wu Chinese, and Hokkien remain scarce. Proprietary LLMs offer commercial service in Cantonese OpenAI et al. (2024); Team et al. (2024) and several Cantonese–Mandarin or Cantonese–English parallel corpora exist in the form of subtitles Wong and Zhang (2017), dictionaries Mair and DeFrancis (2003), or government transcripts Lee (2011). However, resources are limited in scale or accessibility as surveyed by Xiang et al. (2024). As a result, previous work in Cantonese MT has relied on synthetic data augmentation (Liu, 2022; Hong et al., 2024). To the best of our knowledge, the recent addition to FLORES+ Yu et al. (2024) represents the sole publicly available MT resource in Wu Chinese. Hokkien resources comprise of a code-mixing corpus (Lu et al., 2022), a speech-to-speech translation dataset and model (Chen et al., 2023), and orthography standardization for machine translation (Lu et al., 2024a). Beyond Cantonese, Wu, and Hokkien, resources in Hakka (Hung and Huang, 2022; Lai et al., 2024) have been compiled.
$2.$  Language Terminology
We note that we use the terms Cantonese, Wu (Chinese), and Hokkien to describe the Sinitic languages our dataset covers. Some may describe the dataset to cover Yue, Shanghainese, or Southern Min; we acknowledge that names and boundaries between languages in China are often fuzzy (Chappell, 2015), with varying conventions across fields. We clarify that this work discusses annotations in the language rather than dialects specific to the city or a region, although the annotations may reflect the prestige dialect that is spoken in Shanghai, Pearl Delta Region, and Taiwan respectively (Chappell, 2015; Eberhard et al., 2023).
Yue or Cantonese?
The distinction between Yue and Cantonese can be unclear. While Eberhard et al. (2023) describes Cantonese as an alternate name for Yue, some use it to describe the Guangzhou variant of Yue (e.g. Matthews and Yip, 2011). In this work, we follow Ethnologue (Eberhard et al., 2023) and previous work in Cantonese MT (Liu, 2022; Hong et al., 2024) to refer to the entire Yue language as Cantonese. We note that NLLB Team et al. (2024) use "Yue Chinese" to describe what we describe as Cantonese in this paper.
Wu or Shanghainese?
Hokkien or Min Nan?
Hokkien is a variant of Min Nan or Southern Min, spoken primarily in Fujian and Taiwan, as well as in Chinese diasporas in southeast Asia (Eberhard et al., 2023). Previous NLP and language resource work are described as work on Taiwanese Hokkien (Lu et al., 2022; Chen et al., 2023; Lu et al., 2024a; Chu et al., 2025). We note that one exception (Zhang et al., 2025b) uses the term Hokkien. Following the annotators’ preference, we use the term Hokkien.
3.  Dataset annotation
@@
Each example in the dataset consists of a triplet of sentences: a source sentence (src), a machine-translated sentence (mt), and a reference translation (ref). For en-zh, en-yue, and en-wu, the src and ref pairs are drawn from FLORES+, which is an extension of the FLORES-200 dataset (NLLB Team et al., 2024). For Hokkien, the pairs are drawn from The Dictionary from the Ministry of Education [^4] https://sutian.moe.edu.tw/zh-hant/. The mt sentences are generated using the 600M NLLB-200 model Team et al. (2022) for Mandarin and Cantonese, Qwen2.5 Max (Team, 2024) for Wu Chinese, and Taigi-LLaMA-2-Chat-7B Lu et al. (2024b) for Hokkien.
The constructed dataset is then provided to annotators for error span annotation. For each FLORES+ language pair (English-Mandarin, English-Cantonese, English-Wu Chinese), we recruit at least three bilingual annotators. The Hokkien component of SiniticMTError represents a preliminary extension distinct from the other language pairs, and translates from Mandarin to Hokkien.
@@
The annotators are native speakers of their respective target languages and highly proficient in English. They are familiar with both the linguistic conventions of their target language and the goals of the annotation task.
We use the TranslationCorrect tool (Wasti et al., 2025) as our annotation interface.
3.1.  mt Generation
In selecting the model to generate mt sentences, we consider several factors. The ideal mt model would be reproducible, accessible, error-prone, but also reliably able to generate plausible sentences in the target language. We select the 600M NLLB model (Team et al., 2022) for Cantonese and Mandarin, Qwen 2.5 Max (Team, 2024) for Wu, and Taigi-LLaMA-2-Chat-7B (Lu et al., 2024a) for Hokkien. We describe our mt model selection and generation process below.
3.1.1.  Cantonese
For accessibility and reproducibility, we consider open-source models with less than 7B parameters that also have Cantonese proficiency. To verify their reliability in Cantonese generation and error-prone generation, we manually review Cantonese translations of the first 10 English sentences from FLORES+ (Yu et al., 2024) by each model as a preliminary sanity check.
We determine whether the output was in Cantonese and free of language confusion by checking for traditional Han orthography and Cantonese-specific characters. Then, we evaluate mt quality by comparing them to their respective ref sentences, using two metrics: SacreBLEU Post (2018) and ChrF++ Popović (2017).
Out of 600M NLLB-200 (Team et al., 2022), 1.5B Qwen 2.5 Instruct (Team, 2024), 1B Llama 3.2 Instruct (Grattafiori et al., 2024), 1.1B Bloomz (Muennighoff et al., 2023), 1.2B mT0 Large (Muennighoff et al., 2023), 7B Qwen Chat (Bai et al., 2023), 8B Llama 3.1 (Grattafiori et al., 2024), and 8B Aya Expanse (Dang et al., 2024), only NLLB-200, Aya Expanse, and Llama-8B were able to reliably output Cantonese. NLLB-200 and Aya Expanse are described as having been trained on Cantonese data; we were unable to determine whether Llama and Qwen’s training data included Cantonese. Other models do not explicitly report Cantonese data in their training corpora.
The SacreBLEU (Post, 2018) and ChrF++ (Popović, 2017) scores of the three models evaluated were as shown in Table 1. We select NLLB-200 as our model to generate Cantonese mt sentence due to its lowest average SacreBLEU and ChrF++ scores, small size, and easy accessibility.
Model NLLB-200 Aya Expanse Llama
# Params 600M 8B 8B
SacreBLEU 74.5 79.9 82.1
ChrF++ 75.0 72.7 82.0
3.1.2.  Mandarin
For consistency across languages, we used the same model for Mandarin as we did for Cantonese. As discussed in 3.1.1, 600M NLLB (Team et al., 2022) was selected based on a thorough analysis of model quality and parameter size. Using a single model allows us to maintain consistency in prompts and output formatting across languages. Moreover, NLLB shows decent performances in Mandarin, thus making it an ideal choice for producing the mt outputs.
3.1.3.  Wu Chinese
For Wu Chinese, we chose Qwen 2.5 Max (Team, 2024) for producing mt sentences as it remains the only publicly accessible LLM with stable Wu Chinese ability. While DeepSeek is also able to reliably generate Wu output, we are unable to use DeepSeek due to institutional restrictions. We found that other models including Llama and NLLB-200, which were unable to produce usable Wu Chinese output even after prompt engineering.
3.1.4.  Hokkien
To our knowledge, Hokkien is not reliably supported by any of the general multilingual generative models. Machine translation outputs were produced using Taigi-Llama-2-Chat-7B Lu et al. (2024b), a 7-billion-parameter derivative of LLaMA-2 chat model additionally trained on Taiwanese Hokkien data. The model is not a general multilingual model; it was trained primarily on Hokkien, allowing it to reliably generate Hokkien mt sentences.
Error Category Definition
Register Characteristic of text that uses a level of formality higher or lower than required by the specifications or general language conventions.
Table 2: Definitions of 9 error categories used for error annotations.
Severity Level Definition
Table 3: Definitions of severity levels used for error annotations.
3.2.  Annotation Guidelines
@@
Annotators are instructed to examine the mt sentence with reference to both src and ref, and identify any translation errors by highlighting spans directly on the mt sentences. For each error span, annotators categorize the severity and error type, while recording the erroneous span indices in the mt sentence. The label spaces for severity and type are adopted from MQM guidelines (Lommel et al., 2013) and the AfriCOMET framework (Wang et al., 2024), shown in the tables in Section 3.4.
Error types and severity.
We adapted our error severity and category definitions based on MQM guidelines Lommel et al. (2013) and the AfriCOMET framework (Wang et al., 2024), with modifications informed by language-specific characteristics of Sinitic Languages. After multiple rounds of pilot annotation and qualitative analysis as described in Section 3, we refined the labels to better capture common translation issues observed in our data. Our refinement results in 9 error type labels as outlined in Table 2, and 2 severity labels as outlined in Table 3. More detailed information on the guidelines and modifications can be found in Appendix 11.
Granularity.
@@
When identifying errors, the annotators are asked to be as fine-grained as possible. For example, if a sentence contains two words that are each mistranslated, two separate mistranslation error spans should be recorded. If a single text segment contains multiple errors, the leftmost span with highest severity is to be recorded.
Additional information.
@@
One error type is omission, where content in the source sentence is missing from the translation. In such case, annotators are instructed to insert the missing information the post-editing box and highlight the inserted span with the appropriate type (omission), span, and severity labels.
$3.$  Annotation Workflow
@@
Mandarin pilot.
Mandarin annotators had two rounds of pilot studies. In the first round, each annotator completed 50 examples, which were then reviewed by a language lead. The lead provided feedback and held group discussions to ensure consistency, before annotators completed a second round of 50 examples. Qualitative analysis showed clear improvement in annotation consistency.
Cantonese pilot.
Cantonese annotators followed a more granular setup consisting of 4-rounds. The initial round included 50 examples and the others 10 examples each, completing 80 examples in total. Additional rounds of 10 examples were added to ensure sufficient agreement before proceeding.
Main annotation phase.
During the main annotation phase, annotators worked in batches of 50 to 200 sentences. With each batch, we employed a similar iterative process where annotations were reviewed, recurring issues discussed, and relevant guidelines fine-tuned. During such iterative process, guidelines on annotation granularity and categorizing missing punctuation as omission were established.
Quality assurance.
After completing the first round of annotations (main annotator round), annotators proceed to the Quality Assurance (QA) stage. In this phase, each annotator will be assigned sentences previously annotated by another team member. Each participant reviews these annotations and discusses any discrepancies with the original annotator, should there be any, to improve consistency across annotators. We plan to perform inter-annotator agreement analysis in multiple stages.
Wu pilot.
We select a total of 68 sentences from the Wu subset of FLORES+ (Yu et al., 2024). For Wu annotations, we follow a similar setup to that of Cantonese, with a granular setup consisting of 4 rounds. One cycle of annotations was conducted to finish the pilot annotations.
Hokkien pilot.
A total of 154 sentences were collected for this pilot. We use a custom annotation interface [^5] We did not use TranslationCorrect tool for Hokkien. We recruit two annotators, both certified Taiwanese Hokkien teachers and native speakers. The annotation uses a similar MQM-derived schema to Mandarin, Cantonese, and Wu annotation schemes, but with a slightly different set of categories. Nonetheless, our analysis of Hokkien pilot annotations use a mapping between Hokkien categories and those used for Mandarin, Cantonese, and Wu for comparability and consistency across language. We discuss original Hokkien annotation categories in greater detail in Appendix 11.
This pilot differs substantially from the Mandarin, Cantonese, and Wu Chinese components in both methodology and scope. Nonetheless, the adoption of SiniticMTError framework in Hokkien annotations confirms that our framework and data schema can generalize to additional low-resource Sinitic varieties.
3.4.  Dataset Statistics
We report the distribution of error types in each language on 2009 Mandarin sentences and 1402 Cantonese sentences. We also report error type distribution on our pilots–Wu with 68 sentences and Hokkien with 154 sentences. Tables 4, 5, 6, 7, 8 present the distribution of error categories and severity levels in our current annotation data. On average, each Mandarin sentence contains 1.93 annotation spans, while each Cantonese sentence contains 6.38 spans, showing a much higher span density than Mandarin.